What is the energy change when the temperature of 10.9 grams of solid aluminum is decreased from 36.7 degrees celsius to 20.0 degrees celsius
1
Expert's answer
2019-03-18T16:26:55-0400

The specific thermal capacity of Al is "c = 0.902~\frac{\text{J}}{\text{g}\cdot\degree\text{C}};"

"{\Delta}Q = cm{\Delta}t = 0.902~\frac{\text{J}}{\text{g}\cdot\degree\text{C}}~*~10.9~\text{g}~*~(36.7-20.0)~\degree\text{C} \approx 164.19~\text{J}."